OpenStudy (anonymous):

Help! Write ⟨−4,5,1⟩ as a sum of two vectors.

OpenStudy (amistre64):

what 2 numbers add to get -4? add to get 5? or add to get 1?

OpenStudy (anonymous):

-6+2= -4 2+3= 5 2-1= 1

jimthompson5910 (jim_thompson5910):

so your first vector is <-6,2,2> and the second vector is <2,3,-1>. There are infinitely many ways to do this